# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ Appendix E Full Results of Our Proposed Practical Setting
Strategy Channel-Dependent Channel-Independent Models CTF(ours) TimeXer TimeMixer iTrans. CrossGNN TimesNet PatchTST DLinear Metric CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE ETT1 192 0.336 $\pm$ 0.002 0.371 $\pm$ 0.001 0.353 $\pm$ 0.001 0.381 $\pm$ 0.001 0.345 $\pm$ 0.001 0.372 $\pm$ 0.003 0.354 $\pm$ 0.003 0.383 $\pm$ 0.002 0.356 $\pm$ 0.001 0.373 $\pm$ 0.001 0.373 $\pm$ 0.010 0.397 $\pm$ 0.009 0.342 $\pm$ 0.001 0.373 $\pm$ 0.001 0.352 $\pm$ 0.001 0.373 $\pm$ 0.001 336 0.363 $\pm$ 0.001 0.390 $\pm$ 0.000 0.394 $\pm$ 0.001 0.410 $\pm$ 0.002 0.380 $\pm$ 0.002 0.394 $\pm$ 0.002 0.397 $\pm$ 0.001 0.411 $\pm$ 0.002 0.395 $\pm$ 0.001 0.397 $\pm$ 0.001 0.429 $\pm$ 0.011 0.428 $\pm$ 0.005 0.378 $\pm$ 0.002 0.397 $\pm$ 0.001 0.389 $\pm$ 0.001 0.396 $\pm$ 0.001 768 0.427 $\pm$ 0.002 0.430 $\pm$ 0.001 0.449 $\pm$ 0.003 0.443 $\pm$ 0.002 0.456 $\pm$ 0.002 0.437 $\pm$ 0.002 0.472 $\pm$ 0.003 0.454 $\pm$ 0.001 0.463 $\pm$ 0.002 0.438 $\pm$ 0.002 0.533 $\pm$ 0.026 0.483 $\pm$ 0.008 0.451 $\pm$ 0.004 0.440 $\pm$ 0.003 0.454 $\pm$ 0.003 0.438 $\pm$ 0.003 1152 0.467 $\pm$ 0.001 0.454 $\pm$ 0.003 0.484 $\pm$ 0.005 0.462 $\pm$ 0.001 0.498 $\pm$ 0.003 0.464 $\pm$ 0.004 0.514 $\pm$ 0.002 0.479 $\pm$ 0.003 0.507 $\pm$ 0.002 0.464 $\pm$ 0.003 0.535 $\pm$ 0.020 0.488 $\pm$ 0.012 0.485 $\pm$ 0.002 0.460 $\pm$ 0.002 0.502 $\pm$ 0.002 0.471 $\pm$ 0.003 SolarWind 288 0.322 $\pm$ 0.014 0.383 $\pm$ 0.020 0.337 $\pm$ 0.013 0.410 $\pm$ 0.012 0.369 $\pm$ 0.013 0.433 $\pm$ 0.021 0.367 $\pm$ 0.008 0.422 $\pm$ 0.005 0.347 $\pm$ 0.010 0.416 $\pm$ 0.013 0.347 $\pm$ 0.009 0.395 $\pm$ 0.006 0.326 $\pm$ 0.008 0.400 $\pm$ 0.005 0.336 $\pm$ 0.005 0.441 $\pm$ 0.005 576 0.405 $\pm$ 0.007 0.460 $\pm$ 0.010 0.427 $\pm$ 0.008 0.479 $\pm$ 0.010 0.447 $\pm$ 0.013 0.479 $\pm$ 0.013 0.450 $\pm$ 0.005 0.476 $\pm$ 0.004 0.443 $\pm$ 0.010 0.473 $\pm$ 0.010 0.534 $\pm$ 0.038 0.504 $\pm$ 0.026 0.410 $\pm$ 0.011 0.466 $\pm$ 0.006 0.415 $\pm$ 0.003 0.512 $\pm$ 0.003 864 0.417 $\pm$ 0.014 0.461 $\pm$ 0.014 0.426 $\pm$ 0.033 0.467 $\pm$ 0.027 0.489 $\pm$ 0.013 0.507 $\pm$ 0.014 0.488 $\pm$ 0.027 0.499 $\pm$ 0.014 0.511 $\pm$ 0.010 0.512 $\pm$ 0.007 0.474 $\pm$ 0.020 0.489 $\pm$ 0.013 0.447 $\pm$ 0.019 0.491 $\pm$ 0.017 0.457 $\pm$ 0.005 0.545 $\pm$ 0.004 1152 0.456 $\pm$ 0.010 0.488 $\pm$ 0.009 0.489 $\pm$ 0.017 0.514 $\pm$ 0.015 0.492 $\pm$ 0.023 0.523 $\pm$ 0.011 0.547 $\pm$ 0.011 0.527 $\pm$ 0.007 0.573 $\pm$ 0.012 0.547 $\pm$ 0.010 0.553 $\pm$ 0.010 0.526 $\pm$ 0.010 0.464 $\pm$ 0.009 0.494 $\pm$ 0.005 0.470 $\pm$ 0.001 0.557 $\pm$ 0.001 Weather 144 0.179 $\pm$ 0.003 0.225 $\pm$ 0.003 0.205 $\pm$ 0.007 0.245 $\pm$ 0.005 0.205 $\pm$ 0.017 0.240 $\pm$ 0.014 0.232 $\pm$ 0.003 0.271 $\pm$ 0.005 0.200 $\pm$ 0.007 0.248 $\pm$ 0.011 0.258 $\pm$ 0.022 0.288 $\pm$ 0.018 0.198 $\pm$ 0.014 0.236 $\pm$ 0.008 0.207 $\pm$ 0.001 0.261 $\pm$ 0.001 288 0.242 $\pm$ 0.003 0.275 $\pm$ 0.003 0.276 $\pm$ 0.004 0.291 $\pm$ 0.003 0.285 $\pm$ 0.017 0.300 $\pm$ 0.011 0.286 $\pm$ 0.002 0.305 $\pm$ 0.003 0.257 $\pm$ 0.005 0.296 $\pm$ 0.008 0.259 $\pm$ 0.021 0.285 $\pm$ 0.011 0.292 $\pm$ 0.023 0.320 $\pm$ 0.013 0.260 $\pm$ 0.001 0.308 $\pm$ 0.001 576 0.311 $\pm$ 0.002 0.322 $\pm$ 0.001 0.338 $\pm$ 0.003 0.335 $\pm$ 0.002 0.336 $\pm$ 0.016 0.332 $\pm$ 0.013 0.351 $\pm$ 0.001 0.346 $\pm$ 0.001 0.320 $\pm$ 0.006 0.336 $\pm$ 0.007 0.336 $\pm$ 0.007 0.335 $\pm$ 0.007 0.351 $\pm$ 0.016 0.355 $\pm$ 0.008 0.321 $\pm$ 0.001 0.356 $\pm$ 0.002 864 0.343 $\pm$ 0.003 0.347 $\pm$ 0.001 0.367 $\pm$ 0.004 0.360 $\pm$ 0.002 0.354 $\pm$ 0.004 0.353 $\pm$ 0.004 0.380 $\pm$ 0.002 0.368 $\pm$ 0.002 0.361 $\pm$ 0.009 0.372 $\pm$ 0.015 0.392 $\pm$ 0.023 0.380 $\pm$ 0.016 0.367 $\pm$ 0.003 0.359 $\pm$ 0.012 0.356 $\pm$ 0.005 0.395 $\pm$ 0.008 CHS 120 0.104 $\pm$ 0.002 0.062 $\pm$ 0.000 0.125 $\pm$ 0.005 0.067 $\pm$ 0.001 0.105 $\pm$ 0.004 0.061 $\pm$ 0.001 0.124 $\pm$ 0.007 0.072 $\pm$ 0.001 0.105 $\pm$ 0.005 0.070 $\pm$ 0.005 0.155 $\pm$ 0.009 0.079 $\pm$ 0.001 0.127 $\pm$ 0.007 0.066 $\pm$ 0.001 0.114 $\pm$ 0.003 0.106 $\pm$ 0.012 240 0.165 $\pm$ 0.002 0.087 $\pm$ 0.001 0.183 $\pm$ 0.006 0.091 $\pm$ 0.002 0.188 $\pm$ 0.016 0.090 $\pm$ 0.003 0.174 $\pm$ 0.007 0.092 $\pm$ 0.004 0.168 $\pm$ 0.000 0.102 $\pm$ 0.007 0.217 $\pm$ 0.010 0.104 $\pm$ 0.003 0.191 $\pm$ 0.009 0.092 $\pm$ 0.002 0.177 $\pm$ 0.004 0.132 $\pm$ 0.009 480 0.312 $\pm$ 0.004 0.140 $\pm$ 0.001 0.324 $\pm$ 0.004 0.139 $\pm$ 0.000 0.323 $\pm$ 0.029 0.141 $\pm$ 0.002 0.349 $\pm$ 0.008 0.149 $\pm$ 0.003 0.313 $\pm$ 0.001 0.151 $\pm$ 0.001 0.351 $\pm$ 0.016 0.154 $\pm$ 0.002 0.358 $\pm$ 0.008 0.145 $\pm$ 0.003 0.366 $\pm$ 0.027 0.269 $\pm$ 0.035 840 0.548 $\pm$ 0.008 0.213 $\pm$ 0.004 0.556 $\pm$ 0.019 0.216 $\pm$ 0.002 0.582 $\pm$ 0.016 0.216 $\pm$ 0.002 0.585 $\pm$ 0.009 0.217 $\pm$ 0.001 0.560 $\pm$ 0.004 0.239 $\pm$ 0.009 0.612 $\pm$ 0.035 0.229 $\pm$ 0.005 0.586 $\pm$ 0.006 0.226 $\pm$ 0.003 0.693 $\pm$ 0.068 0.435 $\pm$ 0.049

### v2

@@ Appendix E Full Results of Our Proposed Practical Setting
Approach Channel-Dependent Channel-Independent Irregular modeling Model CTF(ours) TimeXer iTrans. CrossGNN TimesNet TimeMixer++ PatchTST DLinear Hi-Patch tPatchGNN Metric CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE ETT1 192 0.335 $\pm$ 0.001 0.370 $\pm$ 0.001 0.354 $\pm$ 0.001 0.381 $\pm$ 0.001 0.356 $\pm$ 0.002 0.384 $\pm$ 0.002 0.355 $\pm$ 0.001 0.373 $\pm$ 0.001 0.388 $\pm$ 0.008 0.407 $\pm$ 0.007 0.363 $\pm$ 0.009 0.389 $\pm$ 0.005 0.343 $\pm$ 0.001 0.374 $\pm$ 0.001 0.352 $\pm$ 0.001 0.373 $\pm$ 0.001 0.396 $\pm$ 0.025 0.422 $\pm$ 0.002 0.413 $\pm$ 0.015 0.440 $\pm$ 0.009 336 0.364 $\pm$ 0.001 0.390 $\pm$ 0.000 0.394 $\pm$ 0.001 0.409 $\pm$ 0.001 0.396 $\pm$ 0.001 0.410 $\pm$ 0.001 0.393 $\pm$ 0.002 0.396 $\pm$ 0.001 0.414 $\pm$ 0.011 0.424 $\pm$ 0.005 0.399 $\pm$ 0.003 0.412 $\pm$ 0.004 0.378 $\pm$ 0.002 0.397 $\pm$ 0.001 0.390 $\pm$ 0.001 0.397 $\pm$ 0.001 0.421 $\pm$ 0.013 0.440 $\pm$ 0.005 0.429 $\pm$ 0.009 0.450 $\pm$ 0.008 768 0.428 $\pm$ 0.001 0.430 $\pm$ 0.000 0.451 $\pm$ 0.002 0.444 $\pm$ 0.001 0.472 $\pm$ 0.003 0.454 $\pm$ 0.001 0.460 $\pm$ 0.002 0.436 $\pm$ 0.001 0.487 $\pm$ 0.015 0.463 $\pm$ 0.008 0.465 $\pm$ 0.010 0.451 $\pm$ 0.004 0.443 $\pm$ 0.004 0.435 $\pm$ 0.002 0.457 $\pm$ 0.002 0.442 $\pm$ 0.003 0.462 $\pm$ 0.009 0.462 $\pm$ 0.004 0.489 $\pm$ 0.005 0.484 $\pm$ 0.004 1152 0.467 $\pm$ 0.001 0.452 $\pm$ 0.002 0.489 $\pm$ 0.002 0.463 $\pm$ 0.001 0.514 $\pm$ 0.002 0.477 $\pm$ 0.002 0.503 $\pm$ 0.003 0.460 $\pm$ 0.002 0.519 $\pm$ 0.017 0.481 $\pm$ 0.010 0.506 $\pm$ 0.012 0.475 $\pm$ 0.006 0.481 $\pm$ 0.002 0.457 $\pm$ 0.002 0.500 $\pm$ 0.002 0.467 $\pm$ 0.003 0.515 $\pm$ 0.021 0.491 $\pm$ 0.012 0.529 $\pm$ 0.006 0.505 $\pm$ 0.004 ETT2 192 0.264 $\pm$ 0.002 0.312 $\pm$ 0.001 0.269 $\pm$ 0.004 0.319 $\pm$ 0.002 0.282 $\pm$ 0.002 0.329 $\pm$ 0.001 0.269 $\pm$ 0.001 0.318 $\pm$ 0.000 0.282 $\pm$ 0.004 0.328 $\pm$ 0.004 0.290 $\pm$ 0.007 0.333 $\pm$ 0.003 0.276 $\pm$ 0.002 0.323 $\pm$ 0.001 0.287 $\pm$ 0.004 0.347 $\pm$ 0.004 0.297 $\pm$ 0.005 0.345 $\pm$ 0.003 0.286 $\pm$ 0.004 0.344 $\pm$ 0.004 336 0.319 $\pm$ 0.002 0.349 $\pm$ 0.001 0.327 $\pm$ 0.004 0.355 $\pm$ 0.002 0.349 $\pm$ 0.002 0.368 $\pm$ 0.001 0.329 $\pm$ 0.001 0.357 $\pm$ 0.002 0.341 $\pm$ 0.006 0.363 $\pm$ 0.003 0.351 $\pm$ 0.009 0.373 $\pm$ 0.006 0.339 $\pm$ 0.005 0.365 $\pm$ 0.003 0.367 $\pm$ 0.010 0.400 $\pm$ 0.008 0.344 $\pm$ 0.005 0.370 $\pm$ 0.003 0.342 $\pm$ 0.007 0.378 $\pm$ 0.007 768 0.436 $\pm$ 0.004 0.416 $\pm$ 0.002 0.439 $\pm$ 0.003 0.420 $\pm$ 0.003 0.451 $\pm$ 0.005 0.426 $\pm$ 0.003 0.449 $\pm$ 0.001 0.421 $\pm$ 0.001 0.466 $\pm$ 0.009 0.433 $\pm$ 0.005 0.446 $\pm$ 0.010 0.427 $\pm$ 0.006 0.450 $\pm$ 0.003 0.431 $\pm$ 0.001 0.527 $\pm$ 0.020 0.490 $\pm$ 0.012 0.453 $\pm$ 0.011 0.430 $\pm$ 0.006 0.450 $\pm$ 0.011 0.434 $\pm$ 0.006 1152 0.488 $\pm$ 0.004 0.454 $\pm$ 0.004 0.485 $\pm$ 0.004 0.454 $\pm$ 0.004 0.501 $\pm$ 0.010 0.459 $\pm$ 0.004 0.498 $\pm$ 0.002 0.456 $\pm$ 0.002 0.509 $\pm$ 0.018 0.461 $\pm$ 0.007 0.496 $\pm$ 0.011 0.465 $\pm$ 0.005 0.494 $\pm$ 0.003 0.465 $\pm$ 0.003 0.641 $\pm$ 0.022 0.550 $\pm$ 0.012 0.494 $\pm$ 0.004 0.460 $\pm$ 0.002 0.495 $\pm$ 0.013 0.468 $\pm$ 0.006 SolarWind 288 0.333 $\pm$ 0.008 0.399 $\pm$ 0.012 0.341 $\pm$ 0.012 0.414 $\pm$ 0.011 0.362 $\pm$ 0.006 0.419 $\pm$ 0.004 0.341 $\pm$ 0.011 0.403 $\pm$ 0.014 0.362 $\pm$ 0.010 0.404 $\pm$ 0.006 0.356 $\pm$ 0.012 0.411 $\pm$ 0.008 0.330 $\pm$ 0.008 0.404 $\pm$ 0.005 0.343 $\pm$ 0.004 0.450 $\pm$ 0.005 0.372 $\pm$ 0.007 0.438 $\pm$ 0.005 0.378 $\pm$ 0.006 0.451 $\pm$ 0.009 576 0.400 $\pm$ 0.005 0.452 $\pm$ 0.006 0.424 $\pm$ 0.007 0.473 $\pm$ 0.008 0.453 $\pm$ 0.004 0.478 $\pm$ 0.003 0.438 $\pm$ 0.010 0.463 $\pm$ 0.010 0.474 $\pm$ 0.036 0.476 $\pm$ 0.021 0.427 $\pm$ 0.010 0.464 $\pm$ 0.008 0.416 $\pm$ 0.010 0.469 $\pm$ 0.006 0.418 $\pm$ 0.003 0.513 $\pm$ 0.003 0.433 $\pm$ 0.022 0.475 $\pm$ 0.018 0.449 $\pm$ 0.017 0.499 $\pm$ 0.019 864 0.428 $\pm$ 0.009 0.472 $\pm$ 0.008 0.452 $\pm$ 0.021 0.487 $\pm$ 0.017 0.511 $\pm$ 0.014 0.512 $\pm$ 0.007 0.508 $\pm$ 0.011 0.505 $\pm$ 0.007 0.500 $\pm$ 0.016 0.497 $\pm$ 0.012 0.457 $\pm$ 0.016 0.493 $\pm$ 0.010 0.452 $\pm$ 0.018 0.495 $\pm$ 0.015 0.453 $\pm$ 0.005 0.542 $\pm$ 0.004 0.447 $\pm$ 0.010 0.485 $\pm$ 0.008 0.469 $\pm$ 0.019 0.508 $\pm$ 0.018 1152 0.453 $\pm$ 0.009 0.484 $\pm$ 0.009 0.478 $\pm$ 0.012 0.503 $\pm$ 0.010 0.552 $\pm$ 0.010 0.531 $\pm$ 0.006 0.573 $\pm$ 0.010 0.543 $\pm$ 0.009 0.547 $\pm$ 0.012 0.519 $\pm$ 0.012 0.475 $\pm$ 0.014 0.504 $\pm$ 0.006 0.471 $\pm$ 0.008 0.498 $\pm$ 0.005 0.470 $\pm$ 0.001 0.557 $\pm$ 0.001 0.471 $\pm$ 0.008 0.494 $\pm$ 0.006 0.494 $\pm$ 0.011 0.514 $\pm$ 0.009 Weather 144 0.185 $\pm$ 0.001 0.231 $\pm$ 0.001 0.223 $\pm$ 0.004 0.255 $\pm$ 0.004 0.235 $\pm$ 0.002 0.272 $\pm$ 0.005 0.193 $\pm$ 0.007 0.247 $\pm$ 0.011 0.209 $\pm$ 0.014 0.248 $\pm$ 0.009 0.178 $\pm$ 0.002 0.225 $\pm$ 0.002 0.175 $\pm$ 0.014 0.258 $\pm$ 0.017 0.207 $\pm$ 0.000 0.262 $\pm$ 0.001 0.218 $\pm$ 0.018 0.263 $\pm$ 0.017 0.233 $\pm$ 0.016 0.272 $\pm$ 0.011 288 0.247 $\pm$ 0.002 0.279 $\pm$ 0.002 0.275 $\pm$ 0.004 0.294 $\pm$ 0.003 0.289 $\pm$ 0.001 0.308 $\pm$ 0.004 0.257 $\pm$ 0.005 0.297 $\pm$ 0.007 0.288 $\pm$ 0.024 0.305 $\pm$ 0.014 0.248 $\pm$ 0.013 0.279 $\pm$ 0.011 0.247 $\pm$ 0.011 0.315 $\pm$ 0.015 0.261 $\pm$ 0.001 0.309 $\pm$ 0.001 0.277 $\pm$ 0.019 0.306 $\pm$ 0.017 0.288 $\pm$ 0.011 0.308 $\pm$ 0.007 576 0.316 $\pm$ 0.001 0.324 $\pm$ 0.001 0.333 $\pm$ 0.003 0.332 $\pm$ 0.002 0.349 $\pm$ 0.001 0.344 $\pm$ 0.001 0.324 $\pm$ 0.006 0.339 $\pm$ 0.006 0.361 $\pm$ 0.013 0.354 $\pm$ 0.010 0.323 $\pm$ 0.003 0.326 $\pm$ 0.004 0.318 $\pm$ 0.006 0.359 $\pm$ 0.009 0.322 $\pm$ 0.001 0.358 $\pm$ 0.002 0.337 $\pm$ 0.012 0.342 $\pm$ 0.011 0.346 $\pm$ 0.011 0.345 $\pm$ 0.006 864 0.352 $\pm$ 0.000 0.351 $\pm$ 0.001 0.368 $\pm$ 0.006 0.359 $\pm$ 0.002 0.380 $\pm$ 0.002 0.369 $\pm$ 0.001 0.367 $\pm$ 0.011 0.378 $\pm$ 0.015 0.396 $\pm$ 0.024 0.381 $\pm$ 0.016 0.354 $\pm$ 0.004 0.355 $\pm$ 0.004 0.359 $\pm$ 0.000 0.393 $\pm$ 0.000 0.356 $\pm$ 0.005 0.394 $\pm$ 0.007 0.372 $\pm$ 0.011 0.369 $\pm$ 0.009 0.383 $\pm$ 0.001 0.372 $\pm$ 0.001 EPA 96 0.641 $\pm$ 0.007 0.525 $\pm$ 0.003 0.651 $\pm$ 0.025 0.509 $\pm$ 0.007 0.657 $\pm$ 0.006 0.519 $\pm$ 0.003 0.725 $\pm$ 0.031 0.558 $\pm$ 0.013 0.720 $\pm$ 0.013 0.553 $\pm$ 0.008 0.705 $\pm$ 0.030 0.551 $\pm$ 0.019 0.624 $\pm$ 0.004 0.495 $\pm$ 0.002 0.921 $\pm$ 0.013 0.657 $\pm$ 0.004 0.673 $\pm$ 0.005 0.570 $\pm$ 0.003 0.663 $\pm$ 0.007 0.563 $\pm$ 0.004 192 0.782 $\pm$ 0.005 0.591 $\pm$ 0.002 0.873 $\pm$ 0.010 0.612 $\pm$ 0.004 0.888 $\pm$ 0.005 0.618 $\pm$ 0.002 0.949 $\pm$ 0.024 0.661 $\pm$ 0.007 0.922 $\pm$ 0.024 0.639 $\pm$ 0.009 0.923 $\pm$ 0.020 0.640 $\pm$ 0.010 0.856 $\pm$ 0.007 0.602 $\pm$ 0.003 1.054 $\pm$ 0.004 0.716 $\pm$ 0.001 0.811 $\pm$ 0.006 0.634 $\pm$ 0.003 0.796 $\pm$ 0.009 0.629 $\pm$ 0.004 288 0.825 $\pm$ 0.006 0.603 $\pm$ 0.001 0.971 $\pm$ 0.008 0.647 $\pm$ 0.003 0.964 $\pm$ 0.005 0.643 $\pm$ 0.001 1.070 $\pm$ 0.052 0.705 $\pm$ 0.014 1.005 $\pm$ 0.007 0.668 $\pm$ 0.004 1.008 $\pm$ 0.017 0.664 $\pm$ 0.005 0.940 $\pm$ 0.007 0.632 $\pm$ 0.002 1.091 $\pm$ 0.009 0.732 $\pm$ 0.002 0.856 $\pm$ 0.006 0.646 $\pm$ 0.003 0.846 $\pm$ 0.016 0.648 $\pm$ 0.005 384 0.858 $\pm$ 0.008 0.623 $\pm$ 0.003 1.047 $\pm$ 0.017 0.676 $\pm$ 0.005 1.019 $\pm$ 0.004 0.666 $\pm$ 0.001 1.171 $\pm$ 0.076 0.735 $\pm$ 0.016 1.103 $\pm$ 0.038 0.712 $\pm$ 0.014 1.087 $\pm$ 0.035 0.694 $\pm$ 0.010 0.995 $\pm$ 0.006 0.659 $\pm$ 0.003 1.121 $\pm$ 0.009 0.748 $\pm$ 0.002 0.894 $\pm$ 0.003 0.667 $\pm$ 0.002 0.900 $\pm$ 0.009 0.672 $\pm$ 0.002 CHS 120 0.103 $\pm$ 0.002 0.062 $\pm$ 0.000 0.122 $\pm$ 0.003 0.067 $\pm$ 0.001 0.132 $\pm$ 0.007 0.072 $\pm$ 0.001 0.108 $\pm$ 0.005 0.065 $\pm$ 0.005 0.142 $\pm$ 0.017 0.077 $\pm$ 0.003 0.116 $\pm$ 0.006 0.067 $\pm$ 0.003 0.129 $\pm$ 0.007 0.065 $\pm$ 0.001 0.113 $\pm$ 0.003 0.098 $\pm$ 0.012 0.123 $\pm$ 0.003 0.067 $\pm$ 0.001 0.106 $\pm$ 0.003 0.061 $\pm$ 0.001 240 0.166 $\pm$ 0.001 0.086 $\pm$ 0.000 0.178 $\pm$ 0.003 0.089 $\pm$ 0.001 0.184 $\pm$ 0.006 0.095 $\pm$ 0.004 0.168 $\pm$ 0.000 0.092 $\pm$ 0.007 0.210 $\pm$ 0.015 0.103 $\pm$ 0.003 0.173 $\pm$ 0.003 0.089 $\pm$ 0.002 0.194 $\pm$ 0.010 0.090 $\pm$ 0.002 0.182 $\pm$ 0.004 0.144 $\pm$ 0.009 0.184 $\pm$ 0.002 0.091 $\pm$ 0.001 0.174 $\pm$ 0.006 0.087 $\pm$ 0.003 480 0.314 $\pm$ 0.003 0.139 $\pm$ 0.000 0.323 $\pm$ 0.004 0.139 $\pm$ 0.000 0.332 $\pm$ 0.009 0.145 $\pm$ 0.002 0.314 $\pm$ 0.000 0.148 $\pm$ 0.007 0.356 $\pm$ 0.017 0.153 $\pm$ 0.004 0.328 $\pm$ 0.010 0.146 $\pm$ 0.005 0.346 $\pm$ 0.008 0.147 $\pm$ 0.003 0.378 $\pm$ 0.028 0.285 $\pm$ 0.033 0.328 $\pm$ 0.007 0.141 $\pm$ 0.002 0.325 $\pm$ 0.006 0.139 $\pm$ 0.002 840 0.555 $\pm$ 0.004 0.216 $\pm$ 0.003 0.570 $\pm$ 0.013 0.217 $\pm$ 0.002 0.573 $\pm$ 0.010 0.215 $\pm$ 0.001 0.553 $\pm$ 0.004 0.232 $\pm$ 0.009 0.613 $\pm$ 0.042 0.225 $\pm$ 0.005 0.566 $\pm$ 0.006 0.215 $\pm$ 0.003 0.591 $\pm$ 0.006 0.226 $\pm$ 0.003 0.731 $\pm$ 0.072 0.466 $\pm$ 0.049 0.571 $\pm$ 0.008 0.216 $\pm$ 0.003 0.574 $\pm$ 0.012 0.213 $\pm$ 0.003